### v1

@@ Abstract
First-order model counting (FOMC) is the problem of counting the number of models of a sentence in first-order logic. Since lifted inference techniques rely on reductions to variants of FOMC, the design of scalable methods for FOMC has attracted attention from both theoreticians and practitioners over the past decade. Recently, a new approach based on first-order knowledge compilation was proposed. This approach, called Crane, instead of simply providing the final count, generates definitions of (possibly recursive) functions that can be evaluated with different arguments to compute the model count for any domain size. However, this approach is not fully automated, as it requires manual evaluation of the constructed functions. The primary contribution of this work is a fully automated compilation algorithm, called Gantry, which transforms the function definitions into C++ code equipped with arbitrary-precision arithmetic. These additions allow the new FOMC algorithm to scale to domain sizes over $500,000$ times larger than the current state of the art, as demonstrated through experimental results.
@@ 1 Introduction
FOMC algorithms are diverse, with approaches ranging from first-order knowledge compilation (FOKC) to cell enumeration [25], local search [14], and Monte Carlo sampling [7]. (See Appendix A for a more detailed comparison of the state-of-the-art FOMC algorithms.) FOKC-based algorithms are particularly prominent, transforming sentences into structured representations such as circuits or graphs. Even when multiple algorithms can solve the same instance, FOKC algorithms are known to find polynomial-time solutions where the polynomial has a lower degree than other approaches [6]. The recently developed ability of a FOKC algorithm to formulate solutions in terms of recursive functions [6] is also noteworthy since the only other proposed alternative is to guess recursive relations [2]. Notable examples of FOKC algorithms include ForcLift [30] and its extension Crane [6].
The Crane algorithm marked a significant step forward, expanding the range of sentences handled by FOMC algorithms. However, it had notable limitations: it required manual evaluation of function definitions to compute model counts and introduced recursive functions without proper base cases, making it more difficult to use. To address these shortcomings, we present Gantry, a fully automated FOMC algorithm that overcomes the constraints of its predecessor. Gantry can handle domain sizes over $500,000$ times larger than those of previous algorithms and simplifies the user experience by automatically managing base cases and compiling function definitions into efficient C++ programs.
The paper is structured as follows. Section 2 covers the necessary preliminaries, including notation, terminology, and background information about (W)FOMC and FOKC. Section 3 describes all aspects of Gantry:
@@
Then, Section 4 presents our experimental results, demonstrating Gantry’s performance compared to other FOMC algorithms. Finally, Section 5 concludes the paper by discussing promising avenues for future work.
@@ 2 Preliminaries
We begin this section by describing some notation that we will use throughout the paper. Then, in Sections 2.1 and 2.2, we introduce the basic terminology of first-order logic and formally define (W)FOMC. Section 2.3 outlines the principles of FOKC, particularly in the context of Crane. Finally, in Section 2.4, we introduce the algebraic terminology used to describe the output of Crane, i.e., functions and equations that define them.
@@
We use $\mathbb{N}_{0}$ to represent the set of non-negative integers. In both algebra and logic, we write $S\sigma$ to denote the application of a substitution $\sigma$ to an expression $S$, where $\sigma=[x_{1}\mapsto y_{1},x_{2}\mapsto y_{2},\dots,x_{n}\mapsto y_{n}]$ signifies the replacement of all instances of $x_{i}$ with $y_{i}$ for all $i=1,\dots,n$. Additionally, for any variable $n$ and $a,b\in\mathbb{N}_{0}$, let $[a\leq n\leq b]\coloneqq\begin{cases}1&\text{if $a\leq n\leq b$}\\ 0&\text{otherwise}\end{cases}$.
@@ 2.1 First-Order Logic
This section reviews the basic concepts of first-order logic used in FOKC algorithms. We focus specifically on the format used internally by ForcLift and its descendants. See Section 2.3.1 for a brief overview of how Gantry transforms an arbitrary sentence in first-order logic into this internal format.
@@ 2.2 First-Order Model Counting
In this section, we will formally define FOMC and its weighted variant. Although this work focuses on FOMC, computing the FOMC using Gantry requires using WFOMC for sentences with existential quantifiers. For such sentences, preprocessing (described in Section 2.3.1) introduces predicates with non-unary weights that must be accounted for to compute the correct model count.
@@
Let $\phi$ be a sentence. For each predicate $P/n$ in $\phi$, let ${(\Delta_{i}^{P})}_{i=1}^{n}$ be a list of the corresponding domains. Let $\sigma$ be a map from the domains of $\phi$ to their interpretations as finite sets, ensuring the sets are pairwise disjoint and contain the corresponding constants from $\phi$. A structure of $\phi$ is a set $M$ of ground literals defined by adding to $M$ either $P(\mathbf{t})$ or $\neg P(\mathbf{t})$ for every predicate $P/n$ in $\phi$ and every $n$-tuple $\mathbf{t}\in\prod_{i=1}^{n}\sigma(\Delta_{i}^{P})$. A structure is a model if it makes $\phi$ valid.
@@
{remark*}
The distinctness of domains is significant in two respects. First, in terms of expressiveness, a clause such as $\forall x\in\Delta\text{. }P(x,x)$ is valid if predicate $P$ operates over two copies of the same domain and invalid otherwise. Second, having more distinct domains makes the problem more decomposable for the FOKC algorithm. With distinct domains, the algorithm can make assumptions or deductions about, e.g., the first domain of predicate $P$ without worrying how (or if) they apply to the second domain.
@@ 2.3.1 Preprocessing
This stage transforms an arbitrary sentence into the format described in Section 2.1, primarily by eliminating existential quantifiers. For example, the first conjunct of Equation 1, i.e.,
@@
where $Z/1$ and $S/1$ are two new predicates with $w^{-}(S)=-1$. One can verify that the WFOMC of Equations 3 and 4 is the same.
@@ 2.3.2 Compilation
After preprocessing, Crane compiles the sentence into the triple $(\mathcal{E},\mathcal{F},\mathcal{D})$, where $\mathcal{E}$ is the set of equations, and $\mathcal{F}$ and $\mathcal{D}$ are auxiliary maps. $\mathcal{F}$ maps function names to sentences. $\mathcal{D}$ maps function names and argument indices to domains. $\mathcal{E}$ can contain any number of functions, one of which (which we will always denote as $f$) represents the solution to the FOMC problem. Computing the FOMC for specific domain sizes involves invoking $f$ with those sizes as inputs. $\mathcal{D}$ records this correspondence between function arguments and domains.
@@ 3 Technical Contributions
Figure 1: Using Gantry to compute the model count of a sentence $\phi$. First, Gantry compiles $\phi$ into a set of equations, which form the basis for creating a C++ program. Executing the program with different command line arguments calculates the model count of $\phi$ for various domain sizes. To accomplish this, the CompileWithBaseCases procedure employs Crane, algebraic simplification techniques (denoted as Simplify), and two other auxiliary procedures.
Figure 1 provides an overview of Gantry’s workflow. We will briefly describe and motivate each procedure before going into more detail in the corresponding subsection.
CompileWithBaseCases (see Section 3.1), the core procedure of Gantry, is responsible for completing the function definitions produced by Crane with the necessary base cases. To do so, it may recursively call itself (and Crane) on other sentences. We prove that the number of such recursive calls is upper bound by the number of domains.
Section 3.1 also describes the Simplify procedure for algebraic simplification. It is crucial for simplifying, e.g., a sum of $n$ terms, only some of which are non-zero. More generally, the equations returned by Crane often benefit from easily detectable algebraic simplifications such as $0$ anything = 0 and $\text{anything}^{0}=1$.
@@
With the help of other procedures outlined above, CompileWithBaseCases returns a set of equations that fully cover the base cases of all recursive functions. While these equations can be intriguing and valuable in their own right, users of FOMC algorithms typically expect a numerical answer. Thus, Section 3.5 describes how Gantry compiles these equations into a C++ program that one can execute with different command-line arguments to compute the model count for various combinations of domain sizes.
@@ 3.2 Identifying a Sufficient Set of Base Cases
In the rest of this section, we will show that the base cases identified by FindBaseCases are sufficient for the algorithm to terminate. [^1] Note that characterising the fine-grained complexity of the solutions found by Gantry or other FOMC algorithms is an emerging area of research. These questions have been partially addressed in previous work [6, 23] and are unrelated to the goals of this section. Let $\mathcal{E}$ denote the equations returned by CompileWithBaseCases.
@@
Given an $n$-ary function $f$ in $\mathcal{E}$ and $\mathbf{x}\in\mathbb{N}_{0}^{n}$, the evaluation of $f(\mathbf{x})$ terminates.
We prove Theorem 3.5 using double induction. First, we apply induction to the number of functions in $\mathcal{E}$. Then, we use induction on the arity of the ‘last’ function in $\mathcal{E}$ according to a topological ordering. Before proving Theorem 3.5, we make a few observations about this and previous [6, 30] work.
@@
Section 3.2 prevents mutual recursion and other cyclic scenarios.
@@ 3.5 Generating C++ Code
In this section, we will describe the final step of Gantry as outlined in Figure 1, i.e., translating the set of equations $\mathcal{E}$ into C++ code. Recall that this step is crucial for the usability of the algorithm; otherwise, function definitions would remain purely mathematical, with no convenient way to compute the model count for particular domain sizes. Once a C++ program is produced, it can be executed with different command-line arguments to determine the model count of the sentence for various domain sizes.
@@ 4 Experimental Evaluation
Our empirical evaluation sought to compare the runtime performance of Gantry with the current state of the art, namely FastWFOMC [^2] https://github.com/jan-toth/FastWFOMC.jl [23, 25] and ForcLift [^3] https://github.com/UCLA-StarAI/Forclift. Our experiments involved two versions of Gantry: Gantry-Greedy and Gantry-BFS. Like its predecessor (see Section 2.3.2), Gantry has two modes for applying compilation rules to sentences: one that uses a greedy search algorithm similar to ForcLift and another that combines greedy and BFS.
@@
We ran each algorithm on each benchmark using domains of size $2^{1},2^{2},2^{3}$, and so on until an algorithm failed to handle a domain size due to a timeout (of one hour) or out-of-memory or out-of-precision errors. While we separately measured compilation and inference time, we primarily focus on total runtime, dominated by the latter. We verified the accuracy of the numerical answers using the corresponding integer sequences in the On-Line Encyclopedia of Integer Sequences [16].

### v2

@@ Abstract
First-order model counting (FOMC) is the problem of counting the number of models of a sentence in first-order logic. Since lifted inference techniques rely on reductions to variants of FOMC, the design of scalable methods for FOMC has attracted attention from both theoreticians and practitioners over the past decade. Recently, a new approach based on first-order knowledge compilation was proposed. This approach, called Crane, instead of simply providing the final count, generates definitions of (possibly recursive) functions that can be evaluated with different arguments to compute the model count for any domain size. However, this approach is not fully automated, as it requires manual evaluation of the constructed functions. The primary contribution of this work is a fully automated compilation algorithm, called Crane2, which transforms the function definitions into C++ code equipped with arbitrary-precision arithmetic. These additions allow the new FOMC algorithm to scale to domain sizes over $500,000$ times larger than the current state of the art, as demonstrated through experimental results.
@@ 1 Introduction
FOMC algorithms are diverse, with approaches ranging from first-order knowledge compilation (FOKC) to cell enumeration [25], local search [14], and Monte Carlo sampling [7]. FOKC-based algorithms are particularly prominent, transforming sentences into structured representations such as circuits or graphs. Even when multiple algorithms can solve the same instance, FOKC algorithms are known to find polynomial-time solutions where the polynomial has a lower degree than other approaches [6]. The recently developed ability of a FOKC algorithm to formulate solutions in terms of recursive functions [6] is also noteworthy since the only other proposed alternative is to guess recursive relations [2]. Notable examples of FOKC algorithms include ForcLift [30] and its extension Crane [6].
The Crane algorithm marked a significant step forward, expanding the range of sentences handled by FOMC algorithms. However, it had notable limitations: it required manual evaluation of function definitions to compute model counts and introduced recursive functions without proper base cases, making it more difficult to use. To address these shortcomings, we present Crane2, a fully automated FOMC algorithm that overcomes the constraints of its predecessor. Crane2 can handle domain sizes over $500,000$ times larger than those of previous algorithms and simplifies the user experience by automatically managing base cases and compiling function definitions into efficient C++ programs.
The paper is structured as follows. Section 2 covers the necessary preliminaries, including notation, terminology, and background information about (W)FOMC and FOKC. Section 3 describes all aspects of Crane2:
@@
Then, Section 4 presents our experimental results, demonstrating Crane2’s performance compared to other FOMC algorithms. Finally, Section 5 concludes the paper by discussing promising avenues for future work.
@@ 2 Preliminaries
We begin this section by describing some notation that we will use throughout the paper. Then, in Sections 2.1 and 2.2, we introduce the basic terminology of first-order logic and formally define (W)FOMC. Section 2.3 provides a brief overview of the different types of first-order logics commonly used in FOMC literature. Section 2.4 outlines the principles of FOKC, particularly in the context of Crane. Finally, in Section 2.5, we introduce the algebraic terminology used to describe the output of Crane, i.e., functions and equations that define them.
@@
We use $\mathbb{N}_{0}$ to represent the set of non-negative integers. In both algebra and logic, we write $S[x_{1}\mapsto y_{1},x_{2}\mapsto y_{2},\dots,x_{n}\mapsto y_{n}]$ to denote the application of a substitution to an expression $S$, where all instances of $x_{i}$ are replaced with $y_{i}$ for all $i=1,\dots,n$. Additionally, for any variable $n$ and $a,b\in\mathbb{N}_{0}$, let $[a\leq n\leq b]\coloneqq\begin{cases}1&\text{if $a\leq n\leq b$}\\ 0&\text{otherwise}\end{cases}$.
@@ 2.1 First-Order Logic
This section reviews the basic concepts of first-order logic used in FOKC algorithms. We focus specifically on the format used internally by ForcLift and its descendants. See Section 2.4.1 for a brief overview of how Crane2 transforms an arbitrary sentence in first-order logic into this internal format.
@@ 2.2 First-Order Model Counting
In this section, we will formally define FOMC and its weighted variant. Although this work focuses on FOMC, computing the FOMC using Crane2 requires using WFOMC for sentences with existential quantifiers. For such sentences, preprocessing (described in Section 2.4.1) introduces predicates with non-unary weights that must be accounted for to compute the correct model count.
@@
Let $\phi$ be a sentence. For each predicate $P/n$ in $\phi$, let ${(\Delta_{i}^{P})}_{i=1}^{n}$ be a list of the corresponding domains. Let $\sigma$ be a map from the domains of $\phi$ to their interpretations as finite sets, ensuring the sets are pairwise disjoint and contain the corresponding constants from $\phi$. A structure of $\phi$ is a set $M$ of ground literals. It is defined by adding to $M$ either $P(\mathbf{t})$ or $\neg P(\mathbf{t})$ for every predicate $P/n$ in $\phi$ and every $n$-tuple $\mathbf{t}\in\prod_{i=1}^{n}\sigma(\Delta_{i}^{P})$. (Here, $\prod$ is the usual Cartesian product.) A structure is a model if it makes $\phi$ valid.
@@ A WFOMC instance comprises:
2.3 The Three Logics of FOMC
Logic Sorts Constants Variables Quantifiers Add. atoms
$\mathsf{FO}$ one or more unlimited $\forall$, $\exists$ $x=y$
$\mathsf{UFO}^{2}+\mathsf{CC}$ $\forall$ $|P|=m$
FOMC commonly utilises three types of first-order logic: $\mathsf{FO}$, $\mathsf{C}^{2}$, and $\mathsf{UFO}^{2}+\mathsf{CC}$. Table 1 summarises the key differences among them. $\mathsf{FO}$ is the input format for ForcLift and its extensions Crane and Crane2. $\mathsf{C}^{2}$ is often used in the literature on FastWFOMC and related methods [11, 12]. (Note that no algorithm accepts $\mathsf{C}^{2}$ as input.) Finally, $\mathsf{UFO}^{2}+\mathsf{CC}$ is the input format supported by the most recent implementation of FastWFOMC [23]. All three logics are function-free, and domains are always assumed to be finite. As usual, we presuppose the unique name assumption, which states that two constants are equal if and only if they are the same constant [18].
In $\mathsf{FO}$, each term has a designated sort, and each predicate $P/n$ corresponds to a sequence of $n$ sorts. Each sort has its corresponding domain. These assignments to sorts are typically left implicit and follow from the quantifiers, e.g., $\forall x,y\in\Delta\text{. }P(x,y)$ implies that the variables $x$ and $y$ have the same sort. On the other hand, $\forall x\in\Delta\text{. }\forall y\in\Gamma\text{. }P(x,y)$ implies that $x$ and $y$ have different sorts, and it would be improper to write, for example, $\forall x\in\Delta\text{. }\forall y\in\Gamma\text{. }P(x,y)\lor x=y$. $\mathsf{FO}$ is also the only logic to support constants, sentences with more than two variables, and the equality predicate. While we do not explicitly refer to sorts in the rest of the paper, the many-sorted nature of $\mathsf{FO}$ is paramount to the algorithms presented therein.
{remark*}
Compared to $\mathsf{FO}$, $\mathsf{C}^{2}$ and $\mathsf{UFO}^{2}+\mathsf{CC}$ lack support for constants, the equality predicate, multiple domains, and sentences with more than two variables. The advantage that $\mathsf{C}^{2}$ brings over $\mathsf{FO}$ is the inclusion of counting quantifiers. That is, alongside $\forall$ and $\exists$, $\mathsf{C}^{2}$ supports $\exists^{=k}$, $\exists^{\leq k}$, and $\exists^{\geq k}$ for any positive integer $k$. For example, $\exists^{=1}x\text{. }\phi(x)$ means that there exists exactly one $x$ such that $\phi(x)$, and $\exists^{\leq 2}x\text{. }\phi(x)$ means that there exist at most two such $x$. $\mathsf{UFO}^{2}+\mathsf{CC}$, on the other hand, does not support any existential quantifiers but instead incorporates (equality) cardinality constraints. For example, $|P|=3$ constrains all models to have precisely three positive literals with the predicate $P$. Except when stated otherwise, we will write all example sentences in $\mathsf{FO}$.
@@ 2.4.1 Preprocessing
This stage transforms an arbitrary sentence into the format described in Section 2.1, primarily by eliminating existential quantifiers via a process called Skolemization [29]. For example, the first conjunct of Equation 1, i.e.,
@@
where $Z/1$ and $S/1$ are two new predicates with $w^{-}(S)=-1$. The idea behind Skolemization is that the output sentence has more models, but for every additional model with a positive weight, there is an equivalent model with a negative weight. Hence, the WFOMC of Equations 3 and 4 is the same. Moreover, note that in this example, preprocessing introduces some redundancy, i.e., the second and third clauses of Equation 4 could be removed. These redundancies are handled at the beginning of the compilation process described in Section 2.4.2.
To illustrate how Skolemization preserves the WFOMC, we let $\Gamma=\{\,a\,\}$ and Δ={ $\Delta=\{\,1,2\,\}$ and assume that $w^{+}(P)=w^{-}(P)=1$. Here, for conciseness, we omit negative literals when writing models. Equation 3 has three models: $\{\,P(a,1)\,\}$, $\{\,P(a,2)\,\}$, and $\{\,P(a,1),P(a,2)\,\}$, all with weight 1, so the WFOMC is 3. Equation 4 has the following models:
  • $\{\,Z(a)\,\}$,
  • $\{\,Z(a),S(a)\,\}$,
  • $\{\,Z(a),S(a),P(a,1)\,\}$,
  • $\{\,Z(a),S(a),P(a,2)\,\}$,
  • $\{\,Z(a),S(a),P(a,1),P(a,2)\,\}$.
The first model has a weight of -1 because $w^{-}(S)=-1$, while all the other models have a weight of 1. Hence, the WFOMC is $4-1=3$.
@@ 2.4.2 Compilation
After preprocessing, Crane compiles the sentence into the triple $(\mathcal{E},\mathcal{F},\mathcal{D})$, where $\mathcal{E}$ is the set of equations, and $\mathcal{F}$ and $\mathcal{D}$ are auxiliary maps. $\mathcal{F}$ maps function names to sentences in such a way that the evaluation of the former corresponds to the FOMC of the latter. $\mathcal{D}$ maps function names and argument indices to domains. The equations in $\mathcal{E}$ can define an arbitrary number of functions, one of which (which we will always denote as $f$) represents the solution to the FOMC problem. Computing the FOMC for specific domain sizes involves invoking $f$ with those sizes as inputs. $\mathcal{D}$ records this correspondence between function arguments and domains.
@@ 3 Technical Contributions
Figure 1: Using Crane2 to compute the model count of a sentence $\phi$. First, Crane2 compiles $\phi$ into a set of equations, which form the basis for creating a C++ program. Executing the program with different command line arguments calculates the model count of $\phi$ for various domain sizes. To accomplish this, the CompileWithBaseCases procedure employs Crane, algebraic simplification techniques (denoted as Simplify), and two other auxiliary procedures.
Figure 1 provides an overview of Crane2’s workflow. We will briefly describe and motivate each procedure before going into more detail in the corresponding subsection.
CompileWithBaseCases (see Section 3.1), the core procedure of Crane2, is responsible for completing the function definitions produced by Crane with the necessary base cases. To do so, it may recursively call itself (and Crane) on other sentences. We prove that the number of such recursive calls is upper bound by the number of domains.
Section 3.1 also describes the Simplify procedure for algebraic simplification. It is crucial for simplifying, e.g., a sum of $n$ terms, only some of which are non-zero. More generally, the equations returned by Crane often benefit from easily detectable algebraic simplifications such as $0$ anything = 0 and $\text{anything}^{0}=1$. Note that the latter includes $10^{0}=$ 1, as 1 is the multiplicative identity.
@@
With the help of other procedures outlined above, CompileWithBaseCases returns a set of equations that fully cover the base cases of all recursive functions. While these equations can be intriguing and valuable in their own right, users of FOMC algorithms typically expect a numerical answer. Thus, Section 3.5 describes how Crane2 compiles these equations into a C++ program that one can execute with different command-line arguments to compute the model count for various combinations of domain sizes.
@@ 3.2 Identifying a Sufficient Set of Base Cases
In the rest of this section, we will show that the base cases identified by FindBaseCases are sufficient for the algorithm to terminate. [^1] Note that characterising the fine-grained complexity of the solutions found by Crane2 or other FOMC algorithms is an emerging area of research. These questions have been partially addressed in previous work [6, 23] and are unrelated to the goals of this section.
@@
Let $\mathcal{E}$ denote the set of equations returned by CompileWithBaseCases, and let $f$ be one of the functions defined in $\mathcal{E}$. The evaluation of $f(\mathbf{x})$ terminates for any tuple $\mathbf{x}$ of non-negative integers (of appropriate length).
We prove Theorem 3.5 using double induction. First, we apply induction to the number of functions in $\mathcal{E}$. Then, we use induction on the arity of the ‘last’ function in $\mathcal{E}$ according to a topological ordering. Before proving Theorem 3.5, we make a few observations that are immediate consequences of this and previous work [6, 30].
@@
Section 3.2 prevents mutual recursion and other cyclic scenarios. It arises from the timeline of FOKC and the fact that, while constructing a graph, the algorithm cannot draw an edge to a vertex that does not yet exist.
@@
Indeed, each $\mathtt{expr}$ is defined by a node type, all of which are finite computations listed in previous work [6, 27, 30].
@@
Now that we have handled the special case when $\mathcal{E}$ defines one function, we are ready to prove the main theorem.
@@ 3.5 Generating C++ Code
In this section, we will describe the final step of Crane2 as outlined in Figure 1, i.e., translating the set of equations $\mathcal{E}$ into C++ code. (Note that the compilation does not rely on any unique features of C++; the equations could easily be compiled into any other programming language.) Recall that this step is crucial for the usability of the algorithm; otherwise, function definitions would remain purely mathematical, with no convenient way to compute the model count for particular domain sizes. Once a C++ program is produced, it can be executed with different command-line arguments to determine the model count of the sentence for various domain sizes. A typical C++ program for one of the benchmarks in Section 4 contains about $80100$ lines of code. Moreover, although we focus on FOMC in this work, support for WFOMC could easily be added by using floating-point numbers instead of integers.
@@ 4 Experimental Evaluation
Our empirical evaluation sought to compare the runtime performance of Crane2 with the current state of the art, namely FastWFOMC [^2] https://github.com/jan-toth/FastWFOMC.jl [23, 25] and ForcLift [^3] https://github.com/UCLA-StarAI/Forclift. Our experiments involved two versions of Crane2: Crane2-Greedy and Crane2-BFS. Like its predecessor (see Section 2.4.2), Crane2 has two modes for applying compilation rules to sentences: one that uses a greedy search algorithm similar to ForcLift and another that combines greedy and BFS.
@@
We ran each algorithm on each benchmark using domains of size $2^{1},2^{2},2^{3}$, and so on until an algorithm failed to handle a domain size due to a timeout (of one hour) or out-of-memory or out-of-precision errors. We verified the accuracy of the numerical answers using the corresponding integer sequences in the On-Line Encyclopedia of Integer Sequences [16].